How do you do a algebra problem like 4x+y=14
-12x+5y=-26

Answer by Sarah900(12)

There are actually multiple ways that you can solve this problem. I'll show you two. The first is substitution.


In substitution you want to isolate one of the variables in either equation. It doesn't really matter which you pick so choose whatever will be easiest. In this case solving for y in the first problem will probably be best. When you do you get...
y=-4x+14

You then substitute that for y in the other equation. This yields
-12x+5(-4x+14)=-26


You can now solve the equation for x
-12x-20x+70=26
-32x=-4496
x=3

You can now take this and substitute it back into either of the equations you had before substituting to find what y is equal to.
y=-4(3)+14
y=2


To check your work replace the numbers you got for x and y into the original equations. If they work then your answers are correct.

The second way of solving the problem is through elimination. In elimination you are going to multiply one or both of the equations by a number in an attempt to get one of the variables to cancel out when you add the two equations together. For example in this problem we could multiply the first equation by 3 so that the x will cancel.

(3)4x+(3)y=14(3)
-12x+5y=-26


12x+3y=42
-12x+5y=-26

Now add the two equations together. Note how 12x and -12x cancel leaving.
8y=16
y=2
2 Can then be substituted back into either of the original equations so that you can solve for x. You can check your work in the same manner you could for substitution.